OpenStudy (anonymous):

The question will be posted below:

OpenStudy (anonymous):

\[-\frac{ 3 }{ 4 } = \frac{ -3 }{ 4 } ?\]

OpenStudy (anonymous):

The question mark isn't a variable. Im asking if the two fractions are equivalent.

OpenStudy (holly00d1248):

im goin with no

OpenStudy (anonymous):

You're wrong -3 divided by 4 = -0.75 -0.75 = \[-\frac{ 3 }{ 4 }\]